OpenStudy (anonymous):

A graph crosses the x–axis at (–1, 0) and (3, 0), and its vertex is at (1, 4). Which quadratic function does it correspond to?

OpenStudy (loser66):

What is the quadratic standard form?

OpenStudy (anonymous):

? A graph crosses the x–axis at (–1, 0) and (3, 0), and its vertex is at (1, 4). Which quadratic function does it correspond to? f(x) = –x2 + 2x – 3 g(x) = x2 + 2x + 3 h(x) = –x2 – 2x – 4 i(x) = –x2 + 2x + 3